OpenStudy (anonymous):

Find the exact value of each expression: cos[tan^-1 (-1) + cos^-1 (-4/5)] I know that tan inverse of -1 is -pi/4 Will FAN and MEDAL for help

OpenStudy (anonymous):

Use the sum and difference identity for cosine \[\cos(u+v)=\cos b \cos v-\sin u \sin v\] and use this triangle to figure out sin (cos^-1 (-4/5)) 

OpenStudy (anonymous):